OpenStudy (anonymous):

if a=2 ; b=-5; c=5 how would you solve the quadratic equation?

hartnn (hartnn):

then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

Use the quadratic foumula \[x_1 =\frac{ -b + \sqrt{b^2-4ac}} {2a}\] \[x_2 =\frac{ -b - \sqrt{b^2-4ac}} {2a}\]

OpenStudy (anonymous):

ya but i got 4+- the square root of -24 / 4 does this make sense?

hartnn (hartnn):

your b=-5, how did u get 4+/- ... ?

OpenStudy (anonymous):

sry the b=-4 and the c=5

OpenStudy (anonymous):

You have imaginary roots. Is that what you're worried about?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Why? Is it a physics problem and you are supposed to have real roots? Because there is nothing wrong with the quadratic formula. Double check your coefficients

hartnn (hartnn):

u actually have imaginary roots for this..

OpenStudy (anonymous):

alright can u tell me where to take it from the imaginary roots?

hartnn (hartnn):

how did u get square root of -24 ?

OpenStudy (anonymous):

because with the coefficients, its (-4)^2-4(2)(5) , in the square root right?

hartnn (hartnn):

lol! b=-5 not -4

OpenStudy (anonymous):

no i i mixed up the coefficents in the original question. i was given a=2;b=-4;c=5

hartnn (hartnn):

ok, \(\huge{x=\frac{4 \pm \sqrt{-24}}{4}}=\frac{4 \pm 2\sqrt{-6}}{4}=\frac{2\pm i\sqrt{6}}{2}\)

OpenStudy (anonymous):

alright thank you very much!